# Semantic Communication-Enabled Wireless Adaptive Panoramic Video Transmission

### v1

@@ IV-A Experimental Environment
The proposed APVST is trained on the panoramic video dataset of 360VDS [20], the dataset of which contains 590 panoramic videos in ERP. The panoramic frames are resized to $512\times 256$ pixels and randomly flipped during training. We evaluate the APVST on the VR scene dataset [21], which contains 208 panoramic videos. For the first frame (I-frame) coding of a GoP, we apply nonlinear transform source-channel coding (NTSCC) [22].
@@
In this experiment, the number of blocks in swin transformerv2 is set to $N_{e}=N_{d}=4$. 8 heads and $8\times 8$ window size are used in MHSA. The quantization sets in Deep JSCC of the primary and motion links are set as $\mathcal{Q}\!=\!\left\{0,2,4,6,8,10,16,20,26,32,20,48,56,64,80,96\right\}$ and $\mathcal{Q}^{\mathrm{ml}}\!=\!\left\{0,2,4,6,8,16,32,48\right\}$, respectively. The balance coefficients are set as $\alpha=1/16$ and $\beta=1/16$. In the model training and testing processes, we set the GoP sizes as $N=7$ and $N=4$, respectively. Adam optimizer is adopted as the optimizer. Other parameters are set as: learning rate is ${10}^{-4}$, training batch size is 8, and testing batch size is 1. The whole APVST model was trained on a single A40 GPU for 4 days.
Figure 5: (a) WS-SSIM performance versus channel bandwidth ratio (CBR), (b) WS-SSIM performance versus SNR, and (c) WS-SSIM and CBR in a GoP over the AWGN channel at SNR=10dB.
IV-B Analysis of Experimental Results
@@
To validate the effectiveness of our proposed APVST scheme in terms of panoramic video transmission, we introduce four schemes for comparison. For semantic communication video transmission schemes, we consider “APVST (w/o WA)” and “DVST” [11], where “APVST (w/o WA)” represents the APVST network without WA module. For traditional video transmission schemes, we consider “H.264 + LDPC” and “H.265 + LDPC”, which employ standard video codecs H.264 and H.265 for source coding and low-density parity check (LDPC) for channel coding. For RGB three-channel panoramic frames, we utilize the channel bandwidth ratio (CBR) [23] to indicate the average ratio of valid symbols transmitted in the channel within a GoP, where CBR can be expressed as $\frac{1}{N}\sum_{t=1}^{N}\frac{k_{t}}{H\times W\times 3}$ and $k_{t}$ represents the dimension of $\mathbf{s}_{t}$.
@@
Fig. 4 shows WS-PSNR achieved by the five schemes varies with CBR. The test SNR is set as 10dB. Owing to the incorporation of the WA module and the latitude adaptive module, our proposed APVST can save approximately 20% and 10% on channel bandwidth cost compared with APVST (w/o WA) and DVST, respectively. Compared with H.264 + LDPC and H.265 + LDPC schemes, the proposed APVST can save bandwidth by about 50% and 40%, respectively. These results show the efficiency of our proposed APVST for panoramic video transmission. However, it can be observed that the WS-PSNR achieved by APVST (w/o WA) and DVST are quite similar at lower CBRs. This indicates that the entropy calculated under the guidance of the entropy model and the latitude adaptive network is smaller, and consequently, the impact of $\mathcal{L}^{\prime}{{}_{t}^{\mathrm{la}}}$ in Equation (9) is also reduced.
@@
Fig. 4 shows WS-PSNR achieved by the five schemes varies with SNR. To ensure fairness, the same CBR is maintained for all schemes. At low SNR levels, semantic transmission significantly outperforms traditional video transmission schemes, which means that the proposed APVST has superior noise immunity properties. However, APVST and DVST achieve the same WS-PSNR at low SNR. This is attributed to the WA module, which primarily learns from the semantic feature map $\mathbf{y}_{t}$ and the weight map $\mathbf{w}$ during the training phase, without incorporating noise-related information, thereby reducing the noise immunity of the network.
@@
Fig. 4 shows WS-PSNR and CBR achieved by the three semantic communication schemes vary with frame number within a GoP. Since the I-frame transmission is without a reference frame, the channel bandwidth ratio is higher compared to the next few frames. The continuous and lossy nature of panoramic video frame transmission leads to the WS-PSNR gradual decrease within a GoP. Compared with DVST and APVST (w/o WA) schemes, the proposed APVST achieves overall gains of about 2.4 dB and 1.1 dB in WS-PSNR under the condition of equal total CBR.
@@
V Conclusion
In this paper, we propose the APVST network, which is based on the Deep JSCC structure and an attention mechanism, aimed at achieving semantic wireless adaptive end-to-end transmission of panoramic video. We propose a latitude adaptive module which is tailored to the specific characteristics of panoramic videos to save bandwidth. To pursue the higher WS-PSNR and WS-SSIM, a WA module is proposed, which works by fusing weights and semantic features. Experimental results demonstrate that the proposed APVST scheme outperforms other video transmission schemes. For future work, we will integrate the proposed APVST scheme with rate-splitting multiple access (RSMA) technology to further enhance the end-to-end transmission efficiency of panoramic videos.

### v2

@@ V-A Experimental Environment
The proposed APVST network is trained on the panoramic video dataset of 360VDS [34], the dataset of which contains 590 panoramic videos in ERP. The panoramic frames are resized to $512\times 256$ pixels and randomly flipped during training. We evaluate the APVST on the VR scene dataset [35], which contains 208 panoramic videos. For the first frame (I-frame) coding of a GoP, we apply nonlinear transform source-channel coding (NTSCC) [36].
@@
In the APVST experiment, the number of blocks in swin transformerv2 is set to $N_{e}=N_{d}=4$. 8 heads and $8\times 8$ window size are used in MHSA. The quantization sets in Deep JSCC of the primary and motion links are set as $\mathcal{Q}\!=\!\left\{0,2,4,6,8,10,16,20,26,32,20,48,56,64,80,96\right\}$ and $\mathcal{Q}^{\text{ml}}\!=\!\left\{0,2,4,6,8,16,32,48\right\}$, respectively. The balance coefficients are set as $\alpha=1/16$ and $\beta=1/16$. In the model training and testing processes, we set the GoP sizes as $N=7$ and $N=4$, respectively. Adam optimizer is adopted as the optimizer. Other parameters are set as: learning rate is ${10}^{-4}$, training batch size is 8, and testing batch size is 1. The whole APVST model was trained on a single A40 GPU for 4 days.
@@
To validate the effectiveness of our proposed APVST scheme in terms of panoramic video transmission, we introduce four schemes for comparison. For semantic communication video transmission schemes, we consider “APVST (w/o WA)” and “DVST” [20], where “APVST (w/o WA)” represents the APVST network without WA module. For traditional video transmission schemes, we consider “H.264 + LDPC” and “H.265 + LDPC”, which employ standard video codecs H.264 and H.265 for source coding and low-density parity check (LDPC) for channel coding.
[…]
Figure 9: Average rewards vs. episode of PPO, TD3, SAC, and DDPG algorithms.
V-B Experimental Analysis
@@
Fig. 7a shows WS-PSNR achieved by these five schemes varies with CBR. The test SNR is set as 10dB. Owing to the incorporation of the WA module and the latitude adaptive module, our proposed APVST can save approximately 20% and 10% on channel bandwidth cost compared with APVST (w/o WA) and DVST, respectively. Compared with H.264 + LDPC and H.265 + LDPC schemes, the proposed APVST can save bandwidth by about 50% and 40%, respectively. These results illustrate the efficiency of our proposed APVST for panoramic video transmission. However, we can see that the WS-PSNR achieved by APVST (w/o WA) and DVST are quite similar under low-CBR conditions. This indicates that the entropy calculated under the guidance of the entropy model and the latitude adaptive network is smaller, and consequently, the impact of $\mathcal{L}^{\prime}{{}_{t}^{\mathrm{la}}}$ in Equation (23) is also reduced.
@@
Fig. 7b shows WS-PSNR achieved by these five schemes varies with SNR. To ensure fairness, the same CBR is maintained for all schemes. At low SNR levels, semantic transmission significantly outperforms traditional video transmission schemes, which means that the proposed APVST has superior noise immunity properties. However, APVST and DVST achieve the same WS-PSNR at low SNR. This is attributed to the WA module, which primarily learns from semantic feature map $\mathbf{y}_{t}$ and weight map $\mathbf{w}$ during the training phase, without incorporating noise-related information, thereby reducing the noise immunity of the network.
@@
Fig. 7c shows WS-PSNR and CBR achieved by the three semantic communication schemes vary with frame number within a GoP. Since the I-frame transmission is without a reference frame, the channel bandwidth ratio is higher compared to the next few frames. The continuous and lossy nature of panoramic video frame transmission leads to the WS-PSNR gradual decrease within a GoP. Compared with DVST and APVST (w/o WA) schemes, the proposed APVST achieves overall gains of about 2.4 dB and 1.1 dB in WS-PSNR under the condition of equal total CBR.
@@
Figure 10: Average rewards vs. episode of RSMA, NOMA, and OFDMA schemes for WS-PSNR and WS-SSIM metrics.
Figure 11: CDF curves for users’ total score of RSMA, NOMA, and OFDMA under different (a) maximum latency requirements, (b) minimum WS-PSNR requirements, and (c) minimum WS-SSIM requirements.
Next, we will analyze the experimental results of PPO for RSMA-enabled transmission semantic streams. To confirm the superiority of the PPO-based algorithm in efficiently transmitting semantic information in panoramic video streams, we conducted a comparative analysis with three other reinforcement learning algorithms: twin delayed DDPG (TD3) [39], soft actor-critic (SAC) [40], and deep deterministic policy gradient (DDPG) [41], as illustrated in Fig. 9. The results demonstrate that the PPO-based algorithm quickly reaches a higher reward level and maintains stability early in the experiments, showcasing outstanding convergence speed and stability. In contrast, although the SAC algorithm experiences significant fluctuations initially, it gradually exhibits sustained performance improvement, eventually stabilizing. The TD3 and DDPG algorithms, however, show slower performance growth and more considerable fluctuations throughout the training process. Therefore, due to its excellent rapid convergence capabilities and high stability, the PPO-based algorithm is confirmed as the optimal choice for addressing this issue.
We also compared the performance of three technologies RSMA, NOMA, and OFDMA in transmitting semantic streams for panoramic videos. The results, depicted in Fig. 10, indicate that our proposed RSMA-enabled APVST scheme significantly outperforms the other two techniques. RSMA not only demonstrates faster convergence speeds and higher average rewards across both WS-PSNR and WS-SSIM metrics, but also maintains superior performance throughout the training process. Specifically, for the WS-PSNR metric, RSMA-enabled APVST scheme achieves performance gains of 13% and 20% over NOMA and OFDMA, respectively. For the WS-SSIM metric, the gains are 6% and 12%. These results underscore the efficiency and stability of RSMA in handling the complex demands of multi-user panoramic video transmission.
To further validate the efficiency of RSMA technology in the transmission of panoramic videos, Fig. 11 shows the cumulative distribution function (CDF) curves of total scores for all users. Overall, the RSMA curves are positioned to the right of those for NOMA and OFDMA, indicating that a greater proportion of users achieve higher total scores under RSMA support. Specifically, Fig. 11a demonstrates that as the maximum latency limit is increased, the number of users meeting this criterion rises, correspondingly shifting the CDF curve to the right. Conversely, reducing the latency limit moves the CDF curve to the left. Similarly, Fig. 11b and  11c illustrate the impact on total scores when adjusting the minimum threshold for immersive experience quality. Increasing the minimum requirement for immersive quality results in fewer users meeting the criterion, shifting the CDF curve to the left. Conversely, lowering this standard shifts the curve to the right. These visuals effectively demonstrate the potential of RSMA technology to enhance user experience quality and reduce latency.
Fig. 12 illustrates the impact of varying the balance coefficient $\kappa$ on the optimization outcomes. Specifically, increasing the weight of the transmission delay fraction (i.e., increasing the $\kappa$ value) caters to scenarios with stringent requirements on transmission delays, while reducing the $\kappa$ value correspondingly decreases the optimized transmission delay score. For scenarios demanding higher immersive experience quality, lowering the $\kappa$ value enhances the weight of the immersive experience quality score, thus user experience can be optimized. However, as observed from (27), the mapping from transmission delay to score is nonlinear. When $T_{\max}$ is low, it becomes challenging for the transmission delay score to reach higher levels, leading to an imbalance between the two scores. For instance, the maximum transmission delay score is 0.69, while the maximum immersive experience quality score approaches 1. As shown in Fig. 12b, increasing $\kappa$ results in a downward trend in the average total score, primarily due to the relatively low transmission delay scores.
Figure 12: (a) Average total score, (b) Average latency score, and (c) Average immersive quality score vs. $\kappa$ of RSMA, NOMA, and OFDMA for WS-PSNR and WS-SSIM metrics.
Figure 13: (a) Average total score, (b) Average latency score, and (c) Average immersive quality score vs. bandwidth of RSMA, NOMA, and OFDMA for WS-PSNR and WS-SSIM metrics.
In Fig. 13, we adjust transmission bandwidth $B$ to evaluate the performance of different technologies in transmitting panoramic video content, which also impacts the maximum transmission power $P_{\max}$ of BS. Fig. 13a demonstrates that the RSMA technology significantly outperforms other technologies in terms of average total scores. Specifically, under the same bandwidth conditions, according to the WSS-PSNR metric, RSMA’s average total score is approximately 11% and 24% higher than NOMA and OFDMA, respectively. For the WSS-SSIM metric, the achieved score by RSMA-enabled APVST scheme is even more impressive, surpassing NOMA and OFDMA by 4% and 10%, respectively. This significant performance enhancement can be attributed to the increased channel capacity and reduced transmission delays that come with increased bandwidth. As shown in Fig. 13b and  13c, the transmission delay scores rapidly increase with increasing bandwidth. However, once the bandwidth reaches a certain level, the rate of improvement in delay scores gradually slows. This indicates that while bandwidth expansion can effectively reduce delays within a certain range, beyond a certain threshold, additional increases in bandwidth contribute less to further reductions in delay. Additionally, changes in bandwidth have a negligible impact on immersive experience quality scores.
VI Conclusion
This paper proposes an adaptive panoramic video semantic transmission framework based on RSMA. The proposed framework, APVST, utilizes a Deep JSCC structure and attention mechanism to adaptively extract and encode semantic features from panoramic frames, effectively enhancing spectral efficiency and saving bandwidth. By integrating an entropy model and a dimension-adaptive module, APVST achieves more precise rate control. Furthermore, the proposed weighted self-attention module enhances the quality of the immersive experience for users. Additionally, the framework takes into account the overlap in the FoV when users watch panoramic videos, utilizing RSMA to effectively split the required panoramic video semantic streams into common and private messages. We formulate this transmission process as a joint optimization problem of latency and immersive experience quality, and exploit the PPO-based algorithm to find the optimal solution, aiming to maximize users’ QoS scores. Extensive simulation results demonstrate that our APVST framework significantly outperforms other semantic and traditional video transmission schemes. Moreover, compared to NOMA and OFDMA, RSMA-enabled APVST scheme achieves higher performance gains in the transmission of panoramic videos, confirming the efficiency and superiority of RSMA in handling high-demand video content. These results not only showcase the application potential of RSMA and semantic communication in immersive communication environments but also provide a solid foundation and valuable references for further research.